# On the rank index of projective curves of almost minimal degree

### v1

@@ 2.2. Projected curves
We now discuss a correspondence between binary forms of degree $d$ and points in a projective space $\mathbb{P}^{d}$. Let $\nu_{d}:\mathbb{P}^{1}\to\mathbb{P}^{d}$ be the $d$-uple embedding (or the $d$th Veronese map). Let $\mathcal{C}_{d}:=\nu_{d}(\mathbb{P}^{1})\subseteq\mathbb{P}^{d}$ denote the image of the Veronese map, which is known as the standard rational normal curve.
Given a point $p\in\mathbb{P}^{d}$, consider the vector space of linear forms on $\mathbb{P}^{d}$ that vanish at $p$. Pulling this space back to $\mathbb{P}^{1}$ via $\nu_{d}$ yields a $d$-dimensional vector space of degree $d$ binary forms, which corresponds to a hyperplane in the space of binary forms of degree $d$. We denote by $f_{p}$ the binary form of degree $d$ (unique up to scalar multiplication) that is orthogonal to this hyperplane with respect to the inner product. In particular, one can see that given a binary form $f_{p}$ in the vector space $\mathbb{K}[s,t]_{d}$, the orthogonal complement of $f_{p}$ with respect to the inner product defined by the differential operator is $((f_{p})^{\perp})_{d}$.
@@
Explicitly, for $p\in\mathbb{P}^{d}$, we may express $p$ as a linear combination of $r\leq d+1$ points on $\mathcal{C}_{d}$, say $p=\sum_{i=1}^{r}c_{i}p_{i}$. Setting $p_{i}=\nu_{d}([\alpha_{i}:\beta_{i}])$, we have
@@
In this correspondence, the torus-fixed points $e_{i}\in\mathbb{P}^{d}$ correspond to the monomials $s^{[d-i]}t^{[i]}\in\mathbb{K}[s,t]_{d}$.
@@
$(f_{p})^{\perp}=(S^{3}T-ST^{3},S^{4}-2S^{2}T^{2}+T^{4})\subset\mathbb{K}[S,T].$
Note that there is a unique degree $6$ form $f_{p}$, up to constant multiples, determined by $(f_{p})^{\perp}$:
@@
  • If $p$ lies on $\mathcal{C}_{d}$, then $\mathcal{C}$ is the curve of minimal degree.
@@
In fact, the graded Betti numbers of $\mathcal{C}$ are determined by the $\mathcal{C}_{d}$-rank of $p$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, see [Par07].
Through the identification between points in $\mathbb{P}^{d}$ and binary forms of degree $d$, we can regard the projection map $\pi_{p}$ as a $\mathbb{K}$-linear map $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow((f_{p})^{\perp})_{d}$ on vector spaces. Indeed, for $f_{p}\in\mathbb{K}[s,t]_{d}$ associated with the point $p\in\mathbb{P}^{d}$, $((f_{p})^{\perp})_{d}$ forms a hyperplane orthogonal to $f_{p}$ with respect to the inner product $\circ$. Indeed, given $p$, let $\{p_{1},\dots,p_{d}\}$ be a basis for $p^{\perp}=\{p^{\prime}\in\mathbb{P}^{d}:p\circ p^{\prime}=0\}$ for any inner product $\circ$ on the space $\mathbb{P}^{d}$. Since the linear projection $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ is the map sending $q=\alpha_{0}p+\sum_{i=1}^{d}\alpha_{i}p_{i}$ to $\sum_{i=1}^{d}\alpha_{i}p_{i}$, the linear projection $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow\mathbb{K}[s,t]_{d}$ sends the binary form $f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}\alpha_{i}f_{i}\in\mathbb{K}[s,t]_{d}$ to $\sum_{i=1}^{d}\alpha_{i}f_{i}$, where $\{f_{1},f_{2},\dots,f_{d}\}$ is the basis for $((f_{p})^{\perp})_{d}$.
Furthermore, by associating a point $p\in\mathbb{P}^{d}$ with a binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and applying apolarity to this form, we can obtain deeper insights into the projected curve $\mathcal{C}$. In detail, for a given point $p\in\mathbb{P}^{d}$, the apolar ideal is given by $(f_{p})^{\perp}=(g_{1},g_{2})\subset\mathbb{K}[s,t]$ with $\deg g_{1}\leq\deg g_{2}$. The $\mathcal{C}_{d}$-rank of the point $p$ equals to $\deg(g_{1})$ which also specifies the type of rational normal surface containing the curve $\pi_{p}(\mathcal{C}_{d})$.
@@ 2.3. Quadratic rank index of projective schemes
For a nondegenerate projective scheme $X\subset\operatorname{\mathbb{P}}^{r}$ defined by quadratic polynomials, we denote by $I(X)$ the homogeneous ideal in the polynomial ring $R=\mathbb{K}[z_{0},z_{1},\ldots,z_{r}]$ with standard grading on variables and let $\{Q_{0},\ldots,Q_{t}\}$ be a basis of $I(X)_{2}$. We say $X$ satisfies property $\operatorname{QR}(k)$ for an integer $k$ if $I(X)$ is generated by a collection of quadratic forms of rank at most $k$. We define the quadratic rank-index of a variety $X$ (defined by quadrics) as the minimal number $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. In this section we discuss how to compute the rank index of $X$ given the basis $\{Q_{0},\ldots,Q_{t}\}$.
A general member $Q$ of $I(X)_{2}$ is of the form
$Q=x_{0}Q_{0}+\cdots+x_{t}Q_{t}\quad\mbox{for some}\quad x_{0},\ldots,x_{t}\in\mathbb{K}.$
We identify $\operatorname{\mathbb{P}}(I(X)_{2})$ with $\operatorname{\mathbb{P}}^{t}$ by sending $[Q]$ to $[x_{0}:\cdots:x_{t}]$. Now, let $M_{Q}$ be the $(r+1)\times(r+1)$ symmetric matrix associated with $Q$. Note that its entries are linear forms in $x_{0},\ldots,x_{t}$. For each $1\leq k\leq r$, we define the locus
$\Phi_{k}(X):=\{[Q]~{}|~{}Q\in I(X)_{2}-\{0\},~{}\mbox{rank}(Q)\leq k\}\subset\operatorname{\mathbb{P}}(I(X)_{2})=\operatorname{\mathbb{P}}^{t}$
of all quadratic polynomials of rank at most $k$ in the projective space $\operatorname{\mathbb{P}}^{t}$. Note that $\Phi_{k}(X)$ is a projective algebraic set in $\operatorname{\mathbb{P}}^{t}$ since it is the zero set of the ideal $I(k+1,M)$ of $\mathbb{K}[x_{0},\ldots,x_{t}]$ generated by all $(k+1)\times(k+1)$ minors of $M$. There is a descending filtration of $\operatorname{\mathbb{P}}^{t}$ associated to $X$:
$\emptyset=\Phi_{1}(X)=\Phi_{2}(X)\subset\Phi_{3}(X)\subset\Phi_{4}(X)\subset\cdots\subset\Phi_{r}(X)\subset\operatorname{\mathbb{P}}^{t}$
@@
[MP23, Proposition 2.1.] Keep the previous notations. Then
$\operatorname{rank-index}(X)=\min\{~{}s\mid\sqrt{I(s+1,M_{Q})}\quad\mbox{has no nonzero linear form.}\}$. }
@@ 2.4. Rank $3$ generators
We review the $Q$-map on a given decomposition of the vector bundle over varieties, which produces rank 3 quadratic generators, introduced in [HLMP21]. Let $L=\mathcal{O}_{X}(1)$ and suppose that $X\subset\operatorname{\mathbb{P}}^{r}$ is the linearly normal embedding of $X$ by $L$. Thus there is a natural isomorphism
$f:H^{0}(X,L)\rightarrow R_{1}$
@@
$Q_{A,B}(s,t,h)=f(s\otimes s\otimes h)f(t\otimes t\otimes h)-f(s\otimes t\otimes h)^{2}.$
This map is well-defined and any rank $3$ quadratic generators of $I(X)$ with the line bundle $L=A^{2}\otimes B$ on $X$ lies on the image of the $Q_{A,B}$ map.
@@ 3. Generators of projected curves
Let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ be the projected curve of degree $d$ in $\mathbb{P}^{d-1}$. In this section, we demonstrate the existence of a parameterized curve that is projectively equivalent to $\mathcal{C}$, and we show that $\mathcal{C}$ satisfies property $\operatorname{QR}(4)$. We note that, since the projected curve is not linearly normal, determining whether the projected curves satisfy property $\operatorname{QR}(k)$ is not straightforward.
@@
Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$. Let $g_{1}$ and $g_{2}$ be binary forms that generate the ideal $(f_{p})^{\perp}$, and let $d_{i}=\deg g_{i}$ for $i=1,2$. Then, the curve $\mathcal{C}$ is projectively equivalent to a curve with the following parametrization:
@@
Given $p\in\mathbb{P}^{d}$, we identify $p$ with the binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and let $\tilde{p}\in\mathbb{K}^{d+1}$ be an affine representative of $p$. We view $\tilde{p}$ as a vector whose initial point is the origin and endpoint is $\tilde{p}$. Note that the orthogonal complement of the vector $\tilde{p}$ corresponds to the vector space $((f_{p})^{\perp})_{d}\subset\mathbb{K}[S,T]_{d}$, and
$((f_{p})^{\perp})_{d}$ $=g_{1}\mathbb{K}[S,T]_{d-d_{1}}+g_{2}\mathbb{K}[S,T]_{d-d_{2}}$
$=\langle g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}},g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}}\rangle,$
where $d_{i}=\deg g_{i}$ for $i=1,2$ with $2\leq d_{1}\leq d_{2}$.
Let $\langle\tilde{p}\rangle^{\perp}=W=\{w\mid w\circ\tilde{p}=0\}$. Then the set $\{\tilde{p},w_{1},\dots,w_{d}\}$ forms a basis of $\mathbb{K}^{d+1}$ for any independent set $\{w_{1},\dots,w_{d}\}$ in $W$. Let $\{v_{0},v_{1},\dots,v_{d}\}$ be the dual basis to $\{\tilde{p},w_{1},\dots,w_{d}\}$ in $(\mathbb{K}^{d+1})^{*}$. Note that $v_{i}\circ\tilde{p}=0$ for $1\leq i\leq d$ and $v_{i}\circ w_{j}=\delta_{ij}$ for $0\leq i\leq d$ and $1\leq j\leq d$.
We choose $\{v_{1},v_{2},\dots,v_{d}\}\subset(\mathbb{K}^{d+1})^{*}$ such that each vector in $\{v_{1},v_{2},\dots,v_{d}\}$ corresponds to the binary forms $\{g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}},g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}}\}\subset\mathbb{K}[S,T]_{d}$, and take $\{w_{1},w_{2},\dots,w_{d}\}\subset\mathbb{K}^{d+1}$ such that $\{w_{1},w_{2},\dots,w_{d}\}$ is dual to $\{v_{1},v_{2},\dots,v_{d}\}$.
Since $\{\tilde{p},w_{1},\dots,w_{d}\}$ forms a basis of $\mathbb{K}^{d+1}$, for any $\tilde{q}\in\mathbb{K}^{d+1}$, we have $\tilde{q}=a_{0}\tilde{p}+\sum_{i=1}^{d}(v_{i}\circ\tilde{q})w_{i}$. Therefore, for $q\in\mathbb{P}^{d}$ whose affine representative is $\tilde{q}\in\mathbb{K}^{d+1}$,
$f_{q}=a_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})f_{w_{1}}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})f_{w_{d}}.$
Then $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ sends $q$ to $[g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}]$.
For $q\in\mathcal{C}_{d}=\{[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\beta^{d}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}$, $q$ is identified with
$f_{q}=(\alpha s+\beta t)^{d}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{d-i}\beta^{i}s^{d-i}t^{i}.$
Then the inner products between the basis elements of $((f_{p})^{\perp})_{d}$ and $f_{q}$ are the evaluations of these basis elements at $(\alpha,\beta)$ by Lemma 2.3. Thus, $\pi_{p}(\mathcal{C})$ is projectively equivalent to the curve
$\{[g_{2}(\alpha,\beta)\alpha^{d-d_{2}}:\dots:g_{2}(\alpha,\beta)\beta^{d-d_{2}}:g_{1}(\alpha,\beta)\alpha^{d-d_{1}}:\dots:g_{1}(\alpha,\beta)\beta^{d-d_{1}}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Due to Theorem 3.1, we can immediately see that $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ is contained in a rational surface scroll $S=S(d-d_{2},d-d_{1})=S(d_{1}-2,d_{2}-2)\subset\mathbb{P}^{d-1}$ as a divisor. This provides a constructive proof of the existence of a surface scroll containing the projected curve. Moreover, we can prove that the ideals of the projected curves can be generated by quadratic forms of rank at most four.
@@
For points $p\in\mathcal{C}_{d}^{2}$, the projected curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, as shown in [Par22]. Therefore, we assume that $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$. Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 4$. By Theorem 3.1, the projected curve $\pi_{p}(\mathcal{C}_{d})$ is projectively equivalent to the curve
@@
where $(f_{p})^{\perp}=(g_{1},g_{2})$ with $d_{1}+d_{2}=d+2$.
Therefore, the curve $\mathcal{C}$ is contained in
$\mathcal{S}$ =$=\left\{[\alpha^{d_{1}-2}\gamma:\dots:\beta^{d_{1}-2}\gamma:\alpha^{d_{2}-2}\eta:\dots:\beta^{d_{2}-2}\eta]\mid[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\right\}$
$=\mathcal{S}(d_{1}-2,d_{2}-2)$
and $\mathcal{C}$ is obtained by the section defined by $g_{1}x-g_{2}y=0$, where $x$ and $y$ are parameters of $\mathcal{S}$. Note that $\mathcal{C}$ is a divisor in $\mathcal{S}$ as the class $H+2F$ where $H$ and $F$ are respectively a hyperplane section and a ruling of $\mathcal{S}$.

### v2

@@ 2.2. Projected curves
We now discuss a correspondence between binary forms of degree $d$ and points in a projective space $\mathbb{P}^{d}$. Let $\nu_{d}:\mathbb{P}^{1}\to\mathbb{P}^{d}$ be the $d$-uple embedding (or the $d$th Veronese map). Let $\mathcal{C}_{d}:=\nu_{d}(\mathbb{P}^{1})\subseteq\mathbb{P}^{d}$ denote the image of the $d$th Veronese map, which is known as the standard rational normal curve.
Given a point $p\in\mathbb{P}^{d}$, consider the vector space of linear forms on $\mathbb{P}^{d}$ that vanish at $p$. Pulling this space back to $\mathbb{P}^{1}$ via $\nu_{d}$ yields a $d$-dimensional vector space of degree $d$ binary forms, which corresponds to a hyperplane in the space of binary forms of degree $d$. We denote by $f_{p}$ the binary form of degree $d$ (unique up to scalar multiple) that is orthogonal to this hyperplane with respect to the inner product. For $p=[p_{0}:p_{1}:\ldots:p_{d}]$, we can write $f_{p}=\sum_{i=0}^{d}p_{i}\binom{d}{i}s^{d-i}t^{i}$, unique up to scalar multiple. In particular, one can see that given a binary form $f_{p}$ in the vector space $\mathbb{K}[s,t]_{d}$, the orthogonal complement of $f_{p}$ with respect to the inner product defined by the differential operator is $((f_{p})^{\perp})_{d}$.
@@
Explicitly, for $p\in\mathbb{P}^{d}$, we may express $p$ as a linear combination of $r\leq d+1$ points $p_{i}$ on $\mathcal{C}_{d}$, say $p=\sum_{i=1}^{r}c_{i}p_{i}$. Setting $p_{i}=\nu_{d}([\alpha_{i}:\beta_{i}])$, we have
@@
In this correspondence, the torus-fixed point $e_{i}:=[\delta_{i0}:\delta_{i1}:\cdots:\delta_{id}]\in\mathbb{P}^{d}$ corresponds to the monomial $s^{[d-i]}t^{[i]}\in\mathbb{K}[s,t]_{d}$ for $i=0,1,\dots,d$.
@@
$(f_{p})^{\perp}=(S^{3}T-ST^{3},S^{4}-S^{2}T^{2}+T^{4})\subset\mathbb{K}[S,T].$
Note that $(f_{p})^{\perp}$ determines a unique degree $6$ form $f_{p}$, up to multiplication by a nonzero scalar in $\mathbb{K}$.
@@
  • If $p$ lies on $\mathcal{C}_{d}$, then $\mathcal{C}$ is a curve of minimal degree.
@@
In fact, the graded Betti numbers of $\mathcal{C}$ are determined by the $\mathcal{C}_{d}$-rank of $p$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, see [14].
Through the identification between points in $\mathbb{P}^{d}$ and binary forms of degree $d$, we can regard the projection map $\pi_{p}$ as a $\mathbb{K}$-linear map $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow((f_{p})^{\perp})_{d}$ on vector spaces. Namely, for $f_{p}\in\mathbb{K}[s,t]_{d}$ associated with the point $p\in\mathbb{P}^{d}$, $((f_{p})^{\perp})_{d}$ forms a hyperplane orthogonal to $f_{p}$ with respect to the inner product $\circ$. Indeed, given $p$, let $\{p_{1},p_{2},\dots,p_{d}\}$ be a basis for $p^{\perp}=\{p^{\prime}\in\mathbb{P}^{d}:p\circ p^{\prime}=0\}$ for any inner product $\circ$ on the space $\mathbb{P}^{d}$. Since the linear projection $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ is the map sending $q=\alpha_{0}p+\sum_{i=1}^{d}\alpha_{i}p_{i}$ to $\sum_{i=1}^{d}\alpha_{i}p_{i}$, the linear projection $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow\mathbb{K}[s,t]_{d}$ sends the binary form $f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}\alpha_{i}f_{i}\in\mathbb{K}[s,t]_{d}$ to $\sum_{i=1}^{d}\alpha_{i}f_{i}$, where $\{f_{1},f_{2},\dots,f_{d}\}$ is the basis for $((f_{p})^{\perp})_{d}$.
Furthermore, by associating a point $p\in\mathbb{P}^{d}$ with the binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and applying apolarity to this form, we can obtain deeper insights into the projected curve $\mathcal{C}$. In detail, for a point $p\in\mathbb{P}^{d}$, the apolar ideal is given by $(f_{p})^{\perp}=(g_{1},g_{2})\subset\mathbb{K}[S,T]$ with $\deg g_{1}\leq\deg g_{2}$. The $\mathcal{C}_{d}$-rank of the point $p$ equals $\deg(g_{1})$ which also specifies the type of rational normal surface containing the curve $\pi_{p}(\mathcal{C}_{d})$.
@@ 2.3. Quadratic rank index of projective schemes
For a nondegenerate projective scheme $X\subset\operatorname{\mathbb{P}}^{r}$ defined by quadratic polynomials, we denote by $I(X)$ the homogeneous vanishing ideal of $X$ in the standard graded polynomial algebra $R=\mathbb{K}[x_{0},x_{1},\ldots,x_{d}]$ and we let $\{Q_{0},Q_{1},\ldots,Q_{t<\binom{d+2}{2}}\}$ be a basis of the $\mathbb{K}$-vector space $I(X)_{2}$. We say $X$ satisfies property $\operatorname{QR}(k)$ for an integer $k$, if $I(X)$ is generated by a collection $\{Q_{0},Q_{1},\ldots,Q_{t}\}$ of quadratic forms of rank at most $k$. We define the quadratic rank-index of a variety $X$ (defined by quadrics) as the minimal number $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. In this section we discuss how to compute the rank index of $X$ given the basis $\{Q_{0},Q_{1},\ldots,Q_{t}\}$.
For each field $\mathbb{F}$ (of characteristic $\neq 2$) and each quadratic polynomial $F\in\mathbb{F}[z_{0},z_{1},\ldots,z_{r}]_{2}$ let $M_{F}\in F^{(r+1)\times(r+1)}$ denote the symmetric matrix associated to $F$, so that $(z_{0},z_{1},\ldots,z_{r})M_{F}(z_{0},z_{1},\ldots,z_{r})^{T}=F$. Each member of $I(X)_{2}$ is of the form
[…]
and satisfies $M_{Q(\underline{a})}=\sum_{i=0}^{t}a_{i}M_{Q_{i}}\in K^{(r+1)\times(r+1)}$. Now, let $A_{0},A_{1},\ldots,A_{t}$ be indeterminates, set $\mathbb{F}:=\mathbb{K}(A_{0},A_{1},\ldots,A_{t})$ and consider the polynomial
$Q:=A_{0}Q_{0}+A_{1}Q_{1}+\cdots+A_{t}Q_{t}\in\sum_{i=0}^{t}A_{i}I(X)_{2}\subseteq\mathbb{F}[z_{0},z_{1},\ldots,z_{r}]_{2}$
which satisfies $M_{Q}=\sum_{i=0}^{t}A_{i}M_{Q_{i}}$. In particular, we have
$M_{Q}\in\mathbb{K}[A_{0},A_{1},\ldots,A_{t}]^{(r+1)\times(r+1)}.$
Moreover, for $1\leq k\leq r$, the form $Q(\underline{a})$ (and hence the matrix $M_{Q(\underline{a})}$) is of rank $\leq r$ if and only if $\underline{a}\in\mathbb{K}^{t+1}$ belongs to the zero-set of the ideal $I(k+1,M_{Q})\subseteq\mathbb{K}[A_{0},A_{1},\ldots,A_{t}]$ generated by all $(k+1)\times(k+1)$ minors of the matrix $M_{Q}$. Now, we identify $\mathbb{P}(I(X)_{2})$ with $\mathbb{P}^{t}$ by sending $Q(\underline{a})]=\mathbb{K}^{*}Q(\underline{a})$ to $a_{0}:a_{1}:\cdots:a_{t}]$, and for $1\leq k\leq r$ we define the locus
$\Phi_{k}(X):=\{[a_{0}:a_{1}:\cdots:a_{t}]\mid\operatorname{rank}(Q(a))\leq k\}\subseteq\mathbb{P}(I(X)_{2})=\mathbb{P}^{t}$
corresponding to all quadratic polynomials of rank at most $k$ in $I(X)$. By the above observation $\Phi_{k}(X)$ is a projective algebraic set in $\mathbb{P}^{t}$. Now, we get a descending filtration of $\mathbb{P}^{t}$ associated to $X$:
$\emptyset=\Phi_{1}(X)=\Phi_{2}(X)\subset\Phi_{3}(X)\subset\Phi_{4}(X)\subset\cdots\subset\Phi_{r}(X)\subset\operatorname{\mathbb{P}}^{t}.$
@@
[13, Proposition 2.1.] Keep the previous notations. Then
$\operatorname{rank-index}(X)=\min\{~s\mid\sqrt{I(s+1,M_{Q})}\quad\mbox{contains no non-zero linear form}\}.$
Definition and Remark 2.6.
For each $k\geq 3$, we define $\delta(X,k)$ as the maximum number of linearly independent quadrics of rank $\leq k$ in $I(X)_{2}$. In particular, $\delta(X,k)$ is one more than the dimension of the linear subspace of $\mathbb{P}^{t}$ spanned by $\Phi_{k}(X)$. Therefore, it holds that
$\delta(X,k)$
$=(t+1)-\left(\text{the maximal}~\sharp~\text{of linearly independent linear forms in}~\sqrt{I(k+1,M_{Q})}\right).$
This invariant is introduced in [11, Definition 6.2] and discussed in the proof of [11, Proposition 6.3].
@@ 2.4. Rank $3$ generators
We review the $Q$-map on a given decomposition of the vector bundle over varieties, which produces rank $3$ quadratic generators, introduced in [8]. Let $X$ be as above, let $L=\mathcal{O}_{X}(1)$ and suppose that $X\subset\operatorname{\mathbb{P}}^{r}$ is the linearly normal embedding of $X$ by $L$. Thus there is a natural isomorphism
$\varphi:H^{0}(X,L)\rightarrow R_{1}$
@@
$Q_{A,B}(s,t,h)=\varphi(s\otimes s\otimes h)\varphi(t\otimes t\otimes h)-\varphi(s\otimes t\otimes h)^{2}.$
This map is well-defined and any rank $3$ quadratic generator of $I(X)$ lies in the image of the $Q_{A,B}$-map.
@@ 3. Generators of projected curves
Let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ be the projected curve of degree $d$ in $\mathbb{P}^{d-1}$. In this section, we demonstrate the existence of a parameterized curve that is projectively equivalent to $\mathcal{C}$, and we show that $\mathcal{C}$ satisfies property $\operatorname{QR}(4)$. We note that, since the projected curve is not linearly normal, determining whether this curve satisfies property $\operatorname{QR}(k)$ is not straightforward.
@@
Let $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ and let $f_{p}\in\mathbb{K}[s,t]$ be the binary form corresponding to $p$. Let $g_{1},g_{2}\in\mathbb{K}[S,T]$ be binary forms that generate the ideal $(f_{p})^{\perp}$, and let $d_{i}=\deg g_{i}$ for $i=1,2$ with $2\leq d_{1}\leq d_{2}$ and $d_{1}+d_{2}=d+2$. Then, the projected curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ is projectively equivalent to a curve with the following parametrization:
@@
Observe that the vector space $V:=\{g\in\mathbb{K}[S,T]_{d}\mid g\circ f_{p}=0\}\subseteq\mathbb{K}[S,T]_{d}$ is given by
$V$ $=g_{1}\mathbb{K}[S,T]_{d-d_{1}}+g_{2}\mathbb{K}[S,T]_{d-d_{2}}$
$=\langle g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}},g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}}\rangle.$
Keep in mind that the $d$ vectors
$v_{1}:=g_{2}S^{d-d_{2}},\;v_{2}:=g_{2}S^{d-d_{2}-1}T,\;\ldots,\;v_{d_{1}-1}:=g_{2}T^{d-d_{2}},$
$v_{d_{1}}:=g_{1}S^{d-d_{1}},\;\ldots,\;v_{d-1}:=g_{1}ST^{d-d_{1}-1},\;v_{d}:=g_{1}T^{d-d_{1}}\in\mathbb{K}[S,T]_{d}$
form a basis of the previously introduced subspace $V\subseteq\mathbb{K}[S,T]_{d}$. We consider $\mathbb{K}[s,t]$ as dual to $\mathbb{K}[S,T]_{d}$ by means of the non-degenerate inner product $\circ$ and choose $w_{1},w_{2},\ldots,w_{d}$ as a system dual to $v_{1},v_{2},\ldots,v_{d}$, so that $v_{i}\circ w_{j}=\delta_{ij}$ for $i,j=1,2,\ldots,d$. Observe that $v_{i}\circ f_{p}=0$ for $i=1,2,\ldots,d$ implies that $f_{p}\notin W:=\sum_{j=1}^{d}\mathbb{K}w_{j}$, so that $\mathbb{K}f_{p}+W=\mathbb{K}[s,t]_{d}$. Hence there is a $v_{0}\in\mathbb{K}[S,T]_{d}$ such that $\{v_{0},v_{1},\ldots,v_{d}\}$ is a basis dual to $\{f_{p},w_{1},\ldots,w_{d}\}$.
Since $\{f_{p},w_{1},w_{2},\ldots,w_{d}\}$ is a basis of $\mathbb{K}[s,t]_{d}$, for any $q\in\mathbb{P}^{d}$, we have a presentation
$f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}(v_{i}\circ f_{q})w_{i}=\alpha_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})w_{1}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})w_{d}.$
If we set $\mathbb{P}^{d-1}:=\mathbb{P}(W)=\mathbb{P}(\mathbb{K}w_{1}+\cdots+\mathbb{K}w_{d})$, the projection map $\pi_{p}:\mathbb{P}^{d}\setminus\{p\}\to\mathbb{P}^{d-1}$ sends
$q$ $=[f_{q}]=[\alpha_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})w_{1}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})w_{d}]$
$=[\alpha_{0}:g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}]\in\mathbb{P}^{d}\setminus\{p\}$
to
$g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}].$
Now, let $q=[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\alpha\beta^{d-1}:\beta^{d}]\in\mathcal{C}_{d}$ with $\alpha:\beta]\in\mathbb{P}^{1}$. Then we have $f_{q}=(\alpha s+\beta t)^{d}$ and Lemma˜2.3 implies that
$g_{2}S^{d-d_{2}}\circ f_{q}=d!\,g_{2}(\alpha,\beta)\alpha^{d-d_{2}},\;\ldots\;,g_{1}T^{d-d_{1}}\circ f_{q}=d!\,g_{1}(\alpha,\beta)\beta^{d-d_{1}}.$
This proves that the curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ has the requested parametrization. ∎
By Theorem˜3.1, we can immediately see that $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ is contained in a rational surface scroll $\mathcal{S}=S(d-d_{2},d-d_{1})=S(d_{1}-2,d_{2}-2)\subset\mathbb{P}^{d-1}$ as a divisor. This provides a constructive proof of the existence of a surface scroll containing the projected curve $\mathcal{C}$. Moreover we can prove that the ideal of this curve can be generated by quadratic forms of rank at most four.
@@
For points $p\in\mathcal{C}_{d}^{2}$, the projected curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, as shown in [16]. Therefore, we assume that $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$. Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 4$. By Theorem˜3.1, the projected curve $\pi_{p}(\mathcal{C}_{d})$ is projectively equivalent to the curve
@@
where $(f_{p})^{\perp}=(g_{1},g_{2})$ with $\deg(g_{i})=d_{i}$, $(i=1,2)$, $d_{1}+d_{2}=d+2$, and $2\leq d_{1}\leq d_{2}$.
Therefore, the curve $\mathcal{C}$ is contained in a rational normal surface scroll
$\mathcal{S}$ $=\left\{[\alpha^{d_{1}-2}\gamma:\dots:\beta^{d_{1}-2}\gamma:\alpha^{d_{2}-2}\eta:\dots:\beta^{d_{2}-2}\eta]\mid[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\right\}$
$=S(d_{1}-2,d_{2}-2)$
and $\mathcal{C}$ is obtained by the section defined by $g_{1}x-g_{2}y=0$, where $x$ and $y$ are local coordinates on $\mathcal{S}$. In fact, $\mathcal{C}$ is a divisor on $\mathcal{S}$ in the class $H+2F$, where $H$ and $F$ are respectively a hyperplane section and a ruling of $\mathcal{S}$. Note that Greek letters are used for parameters in the parametrization of $\mathcal{S}$, whereas $x$ and $y$ are local coordinates on $\mathcal{S}$.